OpenStudy (anonymous):

Who is good at tan, sin, and cos? I really need help quickly. Ill give fan and medal :)

OpenStudy (anonymous):

Based on the figure, which equation is true? (4 points) cos x° = tan y° tan x° = sin y° sin x° = cos y° sin x° = tan y°

OpenStudy (anonymous):

@KendrickLamar2014 could you help me please :)

OpenStudy (anonymous):

If you consider triangle ABC, you can see that sinx, which is the opposite side of the triangle divided by the hypotenuse, is 5/10 = 1/2. If you were to now look at the other triangle, what trig function of angle y would give you 1/2?

OpenStudy (anonymous):

Angle C

OpenStudy (anonymous):

I am really confused on where tan cos and sin are on a triangle so could you explain that? @concentrationalizing

OpenStudy (astrophysics):

All you need to know is what each of the ratios mean, and you can do this easily. 

OpenStudy (astrophysics):

Try to apply this to your question :)

OpenStudy (anonymous):

Sin, cos, tan etc are all in reference to the angle you wish to look at. Im on mobile, so i cant really draw anything. And yeah, astro drew it up, lol. But yeah, when considering angle x, 5 is the side opposite that angle and 10 represents the hypotenuse, the longest side of that triangle. So yhe trig function which uses the opposite and hypotenuse sides is sine, which is why i made thw claim sinx = 5/10 = 1/2

OpenStudy (anonymous):

okay give me a second. Im going to try to figure out the answer based on your picture.

OpenStudy (anonymous):

In the end, youre now using astro's diagram as an example to help you know which trig function is needed for the 2nd triangle. We want some trig function of y to equal 1/2

OpenStudy (anonymous):

I am so lost by both of you. I do not understand

hero (hero):

To help quell the confusion, try finding the lengths of the third side for each right triangle using pythagorean theorem.

OpenStudy (anonymous):

Ill let astro or someone else take over since ill be slow on mobile. Dont want too many voices saying different things.

OpenStudy (anonymous):

is sin x° = tan y° correct?

hero (hero):

Did you find the lengths of the third side for each right triangle?

OpenStudy (anonymous):

No i dont know how to

OpenStudy (anonymous):

i dont know how to find a side using pathagean thereom. My lessons are terrible thats why I do not understand this at all.

hero (hero):

So you've never used \(a^2 + b^2 = c^2\) to solve a right triangle?

OpenStudy (anonymous):

never

OpenStudy (anonymous):

Perhaps the attachment calculated by Mathematica will help out.

OpenStudy (anonymous):

i give up on this I still dont understand it at all. I am just gonna fail it and take it in a real school